# EPiC: Cost-effective Search-based Prompt Engineering of LLMs for Code Generation

### v1

@@
Abstract
Large Language Models (LLMs) have seen increasing use in various software development tasks, especially in code generation. The most advanced recent methods attempt to incorporate feedback from code execution into prompts to help guide LLMs in generating correct code, in an iterative process. While effective, these methods could be costly and time-consuming due to numerous interactions with the LLM and the extensive token usage. To address this issue, we propose an alternative approach named Evolutionary Prompt Engineering for Code (EPiC), which leverages a lightweight evolutionary algorithm to evolve the original prompts toward better ones that produce high-quality code, with minimal interactions with LLM. Our evaluation against state-of-the-art (SOTA) LLM-based code generation models shows that EPiC outperforms all the baselines in terms of cost-effectiveness.
Index Terms:
@@ 1 Introduction
LLMs have been used in many software development activities such as software testing, design, requirement engineering, code generation, maintenance, deployment, etc. [30, 32]. Among these activities, code generation using LLMs has demonstrated significant potential.
@@
In LLM-based code generation, various prompt engineering techniques, including zero-shot [5], in-context learning [33, 34], RAG [35], and task-specific methods [36, 37], have been shown to outperform fine-tuned smaller models. The most advanced prompt engineering methods for code generation employ various agent-based approaches [28]. SOTA methods such as Reflexion [20], Language Agent Tree Search (LATS) [21], AgentCoder [22], LDB [23], and MetaGPT [29] are either planning-based or multi-collaborative agents. While effective, these methods can be costly and time-consuming due to numerous interactions with LLMs which results in extensive token usage, making them less attractive in practical settings. For instance, LATS requires on average 3 minutes to generate the implementation of a function with an average of only 6 lines of code, on the MBPP dataset.
@@
Figure 1: The initial failed prompt (left) and the mutated successful prompt (right)
@@
To tackle this limitation, in this paper, we explicitly focus on the cost of prompt engineering for code and propose a cost-effective approach that leverages a lightweight evolutionary search to optimize prompts for code generation. Our evolutionary-based prompt engineering tool for code generation, namely, EPiC, optimizes prompts for code generation by assessing the generated code against a fitness function, i.e., the pass rate of test cases. Our approach consists of two phases: Initial Evaluation (IE) and Evolutionary Prompt Engineering (EPE). The first phase involves primary code generation using an initial prompt and its evaluation using a set of test cases. If a correct solution is not generated, the process moves to the second (EPE) phase, where an initial population of prompts is generated using the initial prompt. Using each prompt in the initial population, we generate its corresponding code by the LLM under study. We then evaluate the generated code against test cases to calculate its fitness score. Candidate prompts are selected out of the population, using a weighted random selection approach. These candidates are then mutated to form the next generation of prompts. The mutation is carried out using two approaches: one utilizes an LLM guided by a prompt that specifies how to perform the mutation, and the other employs vector embeddings of words to find and replace similar words.
Figure 1 illustrates an example of a mutation applied to a given prompt. In this example, GPT-4o, with the initial prompt on the left, is confused by the word “print” and generates an incorrect print statement. However, the evolved prompt on the right uses the word “publish” which results in the expected return statement, thereby passing the provided test cases.
@@
To evaluate the effectiveness of EPiC, we select two widely used code generation benchmark datasets, i.e., Humaneval [8] and MBPP [24]. We compare EPiC against three SOTA LLM-based code generation tools i.e. Reflexion [20], LATS [21], and LDB [23]. We also implement EPiC leveraging two different LLMs: a large closed-source (GPT4-o) and a smaller open-source one (Magic Coder [11]). In addition, we have empirically investigated the impact of different components of our evolutionary algorithm on the cost and effectiveness of the results.
Overall, our results suggest that the cost and effectiveness of code-generation tools do not always increase proportionally. Specifically, EPiC outperforms all compared SOTA tools (measured by pass@1) by 1% to 3% on HumanEval and 2% to 7% on MBPP with either lower or comparable cost. In addition, EPiC is more effective than LATS (highest performing SOTA) but with 80% less cost and is almost as low cost as Reflexion (the lowest cost approach among SOTA), but with 8% better results.
@@
  1. 1.
    To the best of our knowledge, we are the first to explore the code generation task from the cost perspective.
  2. 2.
@@
  3. 3.
    We have demonstrated the cost-effectiveness of EPiC in code generation compared with baseline methods.
  4. 4.
@@
2 Background and Related Work
In this section, we briefly explain the background of LLMs for code generation and prompt engineering, then report the most related work in the context of prompt engineering of LLM for code.
2.1 LLMs for Code Generation
Models for code generation can be divided into two categories, RNN-based models and transformer-based models [1]. RNN-based models are older models that are outperformed by transformers. Transformer-based code generation models are transformers pre-trained on code-related datasets. For instance, CodeBERT [2] is a transformer model based on BERT [3], specifically pre-trained on the CodeSearchNet dataset [4]. Another example is CodeGPT-2, which builds upon GPT-2 [5] and is pre-trained on the same dataset. Lu et al. [6] evaluated the performance of both GPT-2 and CodeGPT-2 in the code generation task and later, Perez et al. [7], fine-tuned GPT-2 using their dataset for code generation.
Newer versions of LLMs, such as GPT-4, have demonstrated improved performance compared to their predecessors across diverse tasks, including code generation. For instance, Codex [8] is a language model developed by OpenAI based on the GPT architecture, similar to GPT-3. However, it’s specifically designed for code generation and understanding. Codex is trained on a diverse range of publicly available code repositories and strongly focuses on understanding and generating programming code. Another model, CodeLlama [9], is a family of large language models for code based on Llama 2 [10] that is available in 7B, 13B, and 34B parameters. DeepSeek Coder is another state-of-the-art LLM that has been used in software tasks such as Code Generation. This LLM has various sizes, ranging from 1B to 33B. Magicoder [11] is another model that has been fine-tuned on both CodeLlama-7b-Python and deepseek-coder-6.7b.
In this research, we employed two different types of LLMs: GPT-4o, as an example of closed-source large models, and MagicCoder as an example of open-source smaller models. We selected GPT-4o due to its high performance among the closed-source LLMs. MagicCoder was chosen as the most cost-effective small-size open-source model, at the time of designing our experiments.
Figure 2: Diagram of EPiC. On the left, the initial evaluation for the initial prompt assessment is depicted and on the right, the evolutionary process of EPiC is illustrated.
2.2 Prompt Engineering for LLMs
@@
There are multiple categories of prompt engineering, including approaches without training, reasoning and logic, reducing hallucination, and evolutionary-based methods [38]. Zero-shot [5] and few-shot [33] prompting fall under the category of approaches without training. Techniques such as chain-of-thought (CoT) prompting [12], Automatic Chain-of-Thought (Auto-CoT) [39], Self-Consistency [40], and knowledge prompting [13] exemplify reasoning and logic-based methods. To reduce hallucination for prompt engineering, techniques such as Retrieval Augmented Generation (RAG) [35], ReAct Prompting [41], and Chain-of-Knowledge (CoK) Prompting [42] are employed. Evolutionary algorithms are utilized to optimize prompts, as demonstrated by EvoPrompt [16] and PromptBreeder [25]. There are other solutions such as Automated Prompt Engineering [14] and Prompt Optimization with Textual Gradients (ProTeGi) [15].
Zero-shot prompting, as described in [5], eliminates the necessity for extensive training data by employing accurately designed prompts to direct the model toward executing new tasks. The model receives a task description without labeled training data and employs its pre-existing knowledge to generate predictions based on the prompt. Few-shot prompting [33] uses a limited number of input-output examples to help models understand tasks, unlike zero-shot prompting which provides no examples. However, this method requires additional tokens, which can be impractical for longer texts. The selection and composition of examples can also significantly influence model behavior and introduce biases.
CoT [12] improves the reasoning abilities of large language models by incorporating intermediate reasoning steps within prompts. This technique breaks down complex tasks into smaller sub-tasks, mimicking human problem-solving. It significantly enhances performance in arithmetic, commonsense, and symbolic reasoning, especially with larger models. Auto-CoT  [39] is an approach designed to enhance the reasoning capabilities of LLMs by automating the generation of intermediate reasoning steps. It automates the creation of reasoning chains and demonstrations using diversity-based sampling to generate multiple reasoning paths.
Automated prompt engineering approaches use an agent or a similar automated engine to interact with LLM and typically get some feedback to improve the prompt. Zhou et al. [14] proposed APE, a framework designed for the automatic generation and selection of instructions. In this framework, they used an LLM to generate instruction candidates and then search for candidate solutions to maximize a selected score function, treating prompt engineering as an optimization problem. Pryzan et al. [15] proposed Prompt Optimization with Textual Gradients (ProTeGi), which improves prompts for LLMs using a non-parametric approach inspired by numerical gradient descent. It uses natural language gradients from training data to critique and edit prompts, employing a process guided by beam search and bandit selection. This approach improved the efficiency of previous prompt editing techniques across various NLP tasks, including the novel challenge of LLM jailbreak detection.
Evolutionary algorithms have been used in the past for prompt engineering of LLMs in generic NLP tasks. For example, EvoPrompt [16], a framework designed to automate the prompt engineering process for LLMs. EvoPrompt begins with an initial population of prompts and iteratively generates new ones using evolutionary operators such as mutation and crossover, which are performed by LLMs. Another similar approach is PromptBreeder [25], which employs task prompts to condition the LLM and mutation prompts to modify task prompts. The interplay between task and mutation prompts drives iterative improvement in PromptBreeder. Similar approaches for prompt enhancement using LLMs through EAs are also proposed in [26] and [27].
@@
In contrast to existing evolutionary prompt engineering techniques, EPiC is particularly tailored for prompt engineering in coding tasks, with a fitness function defined based on the pass rate of test cases. In addition, it focuses on cost-efficiency throughout the process. To achieve this, it minimizes the calls to external LLMs and implements the mutation operator using local lightweight word embeddings libraries.
2.3 Prompt Engineering of LLMs for Code Generation
Zelikman et al. [17] introduced Parsel, a framework enabling automatic implementation and validation of complex algorithms with code LLMs. Using Parsel, they break down algorithmic tasks into structured descriptions written in natural language, then explore various combinations of function implementations using tests.
@@ 2.3 Prompt Engineering of LLMs for Code Generation
DyLAN [19] is a framework designed to boost the performance of Large Language Model (LLM) agents by assembling them into a dynamic team that adapts to different tasks. Unlike traditional methods with fixed agent sets, DyLAN’s architecture adjusts dynamically to the task query, enhancing its versatility. It employs features like inference-time agent selection and early stopping to improve efficiency and effectiveness. Furthermore, DyLAN incorporates an automatic agent team optimization algorithm based on an unsupervised metric called Agent Importance Score, which selects the most effective agents for each task. Empirical results show DyLAN’s success in tasks like reasoning and code generation, achieving significant improvements over single LLM executions. Reflexion [20] is a reinforcement-based framework for this problem where language agents learn from linguistic feedback rather than weight updates. Agents reflect on task feedback verbally and maintain their own reflective text in memory, which helps them make better decisions in subsequent trials. Reflexion is adaptable to different types and sources of feedback signals and shows significant improvements over baseline agents across various tasks. LATS (Language Agent Tree Search) [21] is another search-based framework that combines LLMs’ abilities in planning, acting, and reasoning. LATS draws inspiration from the Monte Carlo tree search and repurposes LLMs’ strengths as agents, value functions, and optimizers. Crucially, LATS incorporates an environment for external feedback to enable more deliberate and adaptive problem-solving beyond existing techniques’ limitations. AgentCoder [22] utilizes a multi-agent framework with specialized agents: the programmer agent, the test designer agent, and the test executor agent. The programmer agent focuses on code generation and refinement based on feedback from the test executor agent, while the test designer agent generates test cases for the code. The test executor agent runs the code with the test cases and provides feedback to the programmer. This collaborative system aims to improve code generation efficiency, outperforming single-agent models and previous strategies. Zhong et al. [23] introduced Large Language Model Debugger (LDB), a framework that segments programs into basic blocks and tracks intermediate variable values during runtime. LDB enables LLMs to focus on simpler code units, verify their correctness block by block, and pinpoint errors effectively.
@@
In contrast to existing methods and to the best of our knowledge, EPiC is the first search-based prompt engineering method for code generation. It employs a lightweight process to identify the optimal solution in a cost-efficient manner. To achieve this, EPiC utilizes a local embedding function to implement mutation operators on text, to reduce the cost of iterative prompt engineering for code generation. It also guides the search over iterations using the fitness function in Section 4.4, which helpsfindg the most effective prompts.
3 Evolutionary Prompt Engineering for Code
@@
(a) Original prompt and its result
(b) Transformed prompt and its result
Figure 3: An unsuccessful original prompt on the left and the transformed successful version of it on the right. The main issue with the original implementation is the incorrect initialization of the Eulerian number. The orange text in both figures highlights the code and text related to the initialization of the function.
@@
3.1 Initial Evaluation
In Step 1 of the IE phase, the initial prompts for code and test case generation are provided to the LLM (the choice of LLM to be used in our case will be discussed in the Section 4) to generate the corresponding code and test cases. Subsequently, in Step 2, the generated code and test cases are sent to the Testcase Evaluation module. evaluatePrompt in Line 4 of Algorithm 1 is responsible for generating the code and evaluating the test cases. If any test case fails on the code we continue with the EPE phase. If all the test cases pass, we stop and report the generated code as the final answer.
Note that test cases can be provided in various ways. One approach is to use developer-provided test cases for evaluation, while another is to generate test cases using LLMs. The challenge with developer-provided test cases is that typically users do not have test cases before implementation, making the approach impractical for a code generation task. On the other hand, the advantage of using LLM-generated test cases is that it makes the process fully automated. This method is currently used in most SOTA code generation tools [20], [21], [22], [18], and [29], as part of their internal evaluation. Consequently, we also opted to use LLMs for test case generation to ensure a fully automated approach, assuming no developer-provided test cases. Specifically, we employed OpenAI’s GPT-4o model to generate test cases with few-shot examples. To ensure the functional correctness of these test cases, we validated them by parsing their Abstract Syntax Trees (AST). A test case is considered valid if it successfully parses into a syntactically correct AST. This approach is the same approach used in Reflexion [20].
The prompts for code and test case generation which are used in our experiments are provided below. These prompts can be modified if necessary depending on the dataset.
Code Generation Prompt You are a Python developer that implements the correct code based on the function description provided. You are given one or more functions to implement. Don’t delete import statements in the code snippet. Use at most 1000 words.
Test Case Generation Prompt You are an AI coding assistant who can write unique, diverse, and intuitive unit tests for functions given the signature and docstring. Examples:

### v2

# Automated Prompt Engineering for Cost-Effective Code Generation Using Evolutionary Algorithm

@@
Abstract.
Large Language Models (LLMs) have seen increasing use in various software development tasks, especially in code generation. The most advanced recent methods attempt to incorporate feedback from code execution into prompts to help guide LLMs in generating correct code in an iterative process. While effective, these methods could be costly due to numerous interactions with the LLM and extensive token usage. To address this issue, we propose an alternative approach named Evolutionary Prompt Engineering for Code (EPiC), which leverages a lightweight evolutionary algorithm to refine the original prompts into improved versions that generate high-quality code, with minimal interactions with the LLM. Our evaluation against state-of-the-art (SOTA) LLM-based code generation agents shows that EPiC not only achieves up to 6% improvement in pass@k but is also 2–10 times more cost-effective than the baselines.
@@ 1. Introduction
LLMs have been used in many software development activities such as software testing (Dakhel et al., 2024; Guilherme and Vincenzi, 2023), design (Dhar et al., 2024; Kolthoff et al., 2023), requirement engineering (Luo et al., 2022), code generation (Liu et al., 2023b; Shinn et al., 2023; Zhou et al., 2023; Huang et al., 2023b; Zhong et al., 2024), maintenance (Kang et al., 2023), etc. (Fan et al., 2023; Wang et al., 2024). Among these activities, code generation using LLMs has demonstrated significant potential.
@@
In LLM-based code generation, various prompt engineering techniques, including zero-shot (Radford et al., 2019), in-context learning (Brown et al., 2020; Liu et al., 2021b), RAG (Retrieval Augmented Generation) (Lewis et al., 2020), and task-specific methods (He et al., 2022; Carta et al., 2023), have been employed. The most advanced prompt engineering methods for code generation employ various agent-based approaches (Shin et al., 2023). SOTA methods such as Reflexion (Shinn et al., 2023), Language Agent Tree Search (LATS) (Zhou et al., 2023), AgentCoder (Huang et al., 2023b), Large Language Model Debugger (LDB) (Zhong et al., 2024), and MetaGPT (Hong et al., 2023) are either planning-based or multi-collaborative agents. While effective, these methods can be costly due to numerous interactions with LLMs, which result in extensive token usage, making them less attractive in practical settings. For instance, as we found in our study (Section 5.2), on the MBPP dataset, LATS requires on average an additional 234k tokens to find the correct implementation of a function that involves multiple interactions with the LLM.
@@
Figure 1. The initial failed prompt (left) and the evolved successful prompt (right)
One key limitation of these approaches is that the initial prompt fed to the LLM is often suboptimal. If we had an approach that systematically improves the prompt with minimal LLM calls, we could find the optimal prompt. For instance, consider a scenario (Figure 1) where a developer’s original prompt is ambiguous. The LLM may repeatedly generate incorrect code until the user (or an agent) clarifies the instructions in subsequent prompts. Each clarification adds new token usage. By contrast, if we systematically evolve the initial prompt and test it, we can converge on correct code more efficiently.
@@
To accomplish this, we propose Evolutionary Prompt Engineering for Code (EPiC) to refine prompts in a structured and cost-effective way. An Evolutionary Algorithm (EA), in broad terms, maintains a population of candidate solutions (in our case, candidate prompts). It repeatedly evaluates these candidates, selects them based on fitness (test success rate), and mutates them. Our approach consists of two phases: Initial Evaluation (IE) and Evolutionary Prompt Engineering (EPE). The first phase involves primary code generation using an initial prompt and its evaluation using a set of test cases. If a correct solution is not achieved, the process moves to the second (EPE) phase, where an initial population of prompts is augmented using the original prompt. For each prompt in the initial population, LLM generates the corresponding code. The generated code is then evaluated against a set of test cases to determine its fitness score. Candidate prompts are selected from the population using a weighted random selection approach and then mutated to form the next generation of prompts. The mutation is carried out using two approaches: one utilizes an LLM guided by a prompt that specifies how to perform the mutation, and the other employs vector embeddings of words to find and replace similar words.
@@
To evaluate the effectiveness of EPiC, we select three widely used code generation benchmark datasets, i.e., Humaneval+ (Liu et al., 2023a), MBPP+ (Liu et al., 2023a), and BigCodeBench (Zhuo et al., 2024). We implement EPiC using three SOTA LLMs: OpenAI’s o3-mini, Deepseek-v3, and Sonnet 3.7. We also compare EPiC against three SOTA LLM-based agentic code generation tools, i.e., Reflexion (Shinn et al., 2023), LATS (Zhou et al., 2023), and LDB (Zhong et al., 2024).
We use the $pass@k$ metric, as introduced in (Chen et al., 2021), for accuracy (effectiveness), which estimates the probability that at least one of the top k generated code samples is correct. To evaluate the cost-effectiveness, we introduce a new metric called Additional Token Usage per Solved Problem (ATSP in Section  4.3), which quantifies the additional tokens consumed for each extra problem solved relative to a baseline prompting method, where lower values indicate better cost-effectiveness. Our experiments demonstrate that our approach not only consumes fewer tokens per additional solved problem, resulting in lower ATSP scores, but also outperforms the baseline in terms of pass@k by up to 6%. In addition, in terms of ATSP, EPiC is twice as cost-effective as the most cost-effective baseline, Reflexion (Shinn et al., 2023), and almost 10 times more cost-effective than LATS (Zhou et al., 2023).
@@
  1. (1)
    To the best of our knowledge, we are the first to explore the code generation task from the cost-effectiveness perspective.
  2. (2)
@@
  3. (3)
    We have demonstrated the cost-effectiveness of EPiC in code generation compared to baseline methods.
  4. (4)
@@
2. Background and Related Work
In this section, we briefly explain prompt engineering for LLMs in general and report the most related work in the context of prompt engineering of LLM for code. In the end, we discuss evolutionary algorithm in general and how it aligns with EPiC.
Figure 2. Diagram of EPiC. On the left, the initial evaluation for the initial prompt assessment is depicted, and on the right, the evolutionary process of EPiC is illustrated.
2.1. Prompt Engineering for LLMs
@@
There are multiple categories of prompt engineering, including approaches without training, reasoning and logic, reducing hallucination, and evolutionary-based methods (Sahoo et al., 2024). Zero-shot (Radford et al., 2019) and few-shot (Brown et al., 2020) prompting fall under the category of approaches without training. Techniques such as chain-of-thought (CoT) prompting (Wei et al., 2022), Automatic Chain-of-Thought (Auto-CoT) (Zhang et al., 2022), Self-Consistency (Wang et al., 2022), and knowledge prompting (Liu et al., 2021a) exemplify reasoning and logic-based methods. To reduce hallucination for prompt engineering, techniques such as Retrieval Augmented Generation (RAG) (Lewis et al., 2020), ReAct Prompting (Yao et al., 2023), and Chain-of-Knowledge (CoK) Prompting (Li et al., 2023) are employed. Evolutionary algorithms are utilized to optimize prompts, as demonstrated by EvoPrompt (Yang et al., 2023) and PromptBreeder (Fernando et al., 2023). There are other solutions, such as Automated Prompt Engineering (Zhou et al., 2022) and Prompt Optimization with Textual Gradients (ProTeGi) (Pryzant et al., 2023).
Zero-shot prompting (Radford et al., 2019) enables task execution without training data by using well-designed prompts, relying on the model’s pre-existing knowledge. Few-shot prompting (Brown et al., 2020) provides limited examples to guide understanding but requires more tokens, making it less practical for long texts and susceptible to bias from example selection. CoT (Wei et al., 2022) improves LLM reasoning by integrating intermediate steps and breaking down complex tasks to improve arithmetic, common sense, and symbolic reasoning, especially in larger models. Auto-CoT (Zhang et al., 2022) automates reasoning step generation using diversity-based sampling to create multiple reasoning paths. Automated prompt engineering leverages agents to refine prompts based on LLM feedback. Zhou et al. (Zhou et al., 2022) introduced APE, which generates and selects instructions by optimizing a scoring function. Pryzan et al. (Pryzant et al., 2023) proposed ProTeGi, a non-parametric method using natural language gradients and beam search to enhance prompt editing, improving efficiency across NLP tasks. Evolutionary algorithms have been used for LLM prompt engineering in NLP tasks. EvoPrompt (Yang et al., 2023) automates this by iteratively refining prompts using mutation and crossover. Similarly, PromptBreeder (Fernando et al., 2023) employs task and mutation prompts for iterative improvement. Similar approaches for prompt enhancement using LLMs through EAs are also proposed in (Li and Wu, 2023) and (Yang and Li, 2023).
@@
In contrast to existing evolutionary prompt engineering techniques, EPiC is particularly tailored for prompt engineering in coding tasks, with a fitness function defined based on the pass rate of test cases. In addition, it focuses on cost-effectiveness throughout the process. To achieve this, it minimizes the calls to external LLMs and implements the mutation operator using local lightweight word embeddings libraries.
2.2. Prompt Engineering of LLMs for Code Generation
@@
DyLAN (Liu et al., 2023b) enhances LLM performance by dynamically assembling agent teams based on task queries, unlike fixed-agent methods. It improves efficiency with inference-time agent selection, early stopping, and an unsupervised Agent Importance Score for optimizing agent selection. Reflexion (Shinn et al., 2023) is a reinforcement-based framework in which language agents learn from linguistic feedback. By reflecting on task feedback and storing reflective text in memory, agents improve decision-making in future trials. LATS (Zhou et al., 2023) (Language Agent Tree Search) is a framework that integrates the planning, acting, and reasoning abilities, inspired by Monte Carlo tree search. It utilizes LLMs as agents, value functions, and optimizers while incorporating external feedback for more adaptive problem-solving. AgentCoder (Huang et al., 2023b) employs a multi-agent framework with specialized roles: the programmer generates and refines code, the test designer creates test cases, and the test executor runs tests and provides feedback. This collaborative approach enhances code generation efficiency, surpassing single-agent models. Zhong et al. (Zhong et al., 2024) introduced Large Language Model Debugger (LDB), a framework that segments programs into basic blocks and tracks intermediate variable values during runtime. LDB enables LLMs to focus on simpler code units and verify their correctness block by block.
@@
In contrast to existing methods and to the best of our knowledge, EPiC is the first evolutionary-based prompt engineering method for code generation. It employs a lightweight process to identify the optimal solution in a cost-effective manner. To achieve this, EPiC utilizes a local embedding function to implement mutation operators on text to reduce the cost of iterative prompt engineering for code generation. It also guides the search over iterations using the fitness function in Section 4.4, which helps in finding the best prompts.
2.3. Evolutionary Algorithms
Evolutionary algorithms (EA) are population-based metaheuristics inspired by natural selection. One key advantage of EA is that it can optimize using only a fitness function without requiring gradient information, making it particularly suitable for problems where gradients are difficult or impossible to compute, such as prompt optimization, where the pass rate of generated code is a discrete and non-differentiable function. EA typically follows a cycle of initialization, evaluation, selection, variation, and iteration. The process begins with an initial population of candidate solutions, which, in EPiC, are variations of the original prompt. Each candidate is evaluated using a fitness function that measures quality based on the pass rate of test cases. Selection then occurs through a weighted random strategy that prioritizes prompts demonstrating higher pass rate. Variation operators, such as mutation, introduce modifications to candidates, and this iterative cycle continues until convergence or a predefined stopping criterion is met.
While EA is effective in navigating complex search spaces, it also has drawbacks, such as potential inefficiency when fitness evaluations are expensive and the risk of premature convergence. However, for optimizing prompts—where evaluating fitness is relatively straightforward but computing gradient information is non-trivial—EA is a natural and effective choice. Over time, this process refines the prompt by focusing on those that lead to code passing more test cases, ultimately producing an optimized prompt whose generated code passes all test cases.
3. Evolutionary Prompt Engineering for Code
@@
3.1. Initial Evaluation
Initial evaluation (IE) is intended for an early stop in the code generation process if a correct solution is identified. In Step 1 of the IE phase, the initial prompts for code and test case generation are provided to the LLM (the choice of LLM to be used in our case will be discussed in the Section 4) to generate the corresponding code and test cases. In Step 2, the generated code is evaluated against the generated test cases. evaluatePrompt in Line 4 of Algorithm 1 is responsible for generating the code and evaluating the test cases. If any test case fails on the code, we continue with the EPE phase. If all test cases pass, we report the generated code as the final answer.
Note that test cases can be provided in various ways. One approach is to use developer-provided test cases for evaluation, while another is to generate test cases using the LLM. The challenge with developer-provided test cases is that, typically, users do not have test cases before implementation, making the approach impractical for a code generation task. On the other hand, the advantage of using LLM-generated test cases is that it makes the process fully automated. This method is currently used in most SOTA code generation tools (Shinn et al., 2023), (Zhou et al., 2023), (Huang et al., 2023b), (Huang et al., 2023a), and (Hong et al., 2023), as part of their internal evaluation. Consequently, we also opted to use LLMs for test case generation to ensure a fully automated approach, assuming no developer-provided test cases. To ensure the functional correctness of these test cases, we validated them by parsing their Abstract Syntax Trees (AST). A test case is considered valid if it successfully parses into a syntactically correct AST. This approach is the same approach used in Reflexion (Shinn et al., 2023).
The prompts for code and test case generation that are used in our experiments are provided below. These prompts can be modified if necessary, depending on the dataset.
Test Case Generation Prompt You are an AI coding assistant who can write unique, diverse, and intuitive unit tests for functions, given the signature and docstring. Use step-by-step reasoning to validate each test case separately to ensure that all of them have the correct expected output. Write the final validated tests inside \$\$\$\$ tags. Examples: